The circumference of a circle is 20Pi. What is the area of the circle? Leave your answer in terms of Pi
Found 2 solutions by stanbon, checkley75:
Answer by stanbon(75887)

The circumference of a circle is 20Pi. What is the area of the circle? Leave your answer in terms of Pi
-----
Circmference = (pi)d = 20pi
Therefore diameter = 20 and radius=10
-------------
Then Area = (pi)r^2 = (pi)10^2 = 100pi
===============
Cheers,
Stan H.

Answer by checkley75(3666)

C=2piR
20pi=2piR DIVIDE BOTH SIDES BY pi
20=2R
R=20/2
R=10 THUS THE AREA OF THIS CIRCLE IS
A=piR^2
A=pi10^2
A=100pi